# GP+: A Python Library for Kernel-based learning via Gaussian Processes

### v1

@@ 5.2 Multi-fidelity Modeling
We use two analytic (Sinusoidal and Wing) and one engineering (DNS-ROM) examples for the comparison (see Table 4 and [105] for details on these examples). Sinusoidal is a $1$-dimensional, bi-fidelity example for which we generate a dataset consisting of $400$ HF and $2000$ LF samples. Wing has $4$ fidelity sources ($1$ HF and $3$ LFs) and we produce $1500$ samples from the HF source and $4000$ samples from each of the LF sources. In both analytic examples, $1\%$ of data is used for training and the rest of the HF data for testing. DNS-ROM is a $5$-dimensional problem on fracture modeling of metallic alloys where the data are generated via four different simulators with $70,110,170,250$ samples. In this example, we use $20\%$ of the samples for training and the rest of the HF data for testing.
The results for each approach on each problem are summarized in Table 2 and demonstrate that GP+ significantly outperforms the other baselines in all problems. More specifically, while V-GP is limited to the small HF data, GP+ effectively leverages the information provided by the LF data to learn the HF source. The poor performance of STMF-GP is due to the fact that from a methodological standpoint it models the inter-relations between the data sources incorrectly. In addition to providing low accuracy, the predictions of STMF-GP are sensitive to the values assigned to its fidelity indices. This is evident in Table 2 where the prediction errors for two different yet close sets of random indices ($STMF-GP_{1}$ and $STMF-GP_{2}$, see also Table 11) are very different. Regarding FFNNs, we attribute their poor performance in all problems to their architecture and, in particular, their simple mechanism for handling fidelity levels and. These FFNNs simply one-hot encode the fidelity indices and ignore the intricate correlations among the corresponding data sources. Compared to other methods, the reported NRMSEs for FFNNs are more sensitive to the model architecture and notably change as the network size varies. This sensitivity is partly due to the small size of the MF data and can perhaps be improved by iteratively refining the architecture or the optimization parameters (e.g., learning rate schedule or regularization weights). However, we avoid such detailed refinements since none of the other methods are fine-tuned.
@@
Wing and DNS-ROM are relatively complex problems with small amounts of data and different types of noise (e.g., inDNS-ROM the noise variance depends on the source while in Wing it does not). GP+ is very well suited to tackle these types of problems because the number of its hyperparameters scales much better than FFNNs, is not limited to the small HF data, and better estimates noise as explained in Section 4.1.1. Accordingly, we observe lower prediction errors for GP+ in these examples.
@@ 5.2.1 Deterministic and Probabilistic Embedding
Figure 9: Probabilistic vs deterministic embedding for MF modeling: While both embeddings estimate the same degree of similarity among sources, the probabilistic one more accurately quantifies uncertainties.
@@
Figure 10: Prediction performance with probabilistic vs deterministic embedding: We report NRMSE and NIS values calculated across $10$ repetitions. Both approaches are very accurate but the probabilistic one is more robust for HF emulation especially since HF data is very small. The similar NRMS and NIS values for LF1 and LF2 are expected because of their high correlation.
Figure 9 illustrates the fidelity embeddings learned through probabilistic and deterministic MF modeling. We only show the learnt embeddings in one of the repetitions since the relative distances across the $10$ repetitions are quite similar (albiet the exact locations are different). As explained in Section 4, these embeddings indicate how similar or correlated different data sources are with respect to each other. Specifically, the latent distance between the points encoding the LF sources from the point that encodes the HF source are consistent with the NRMSE values reported in Table 4 where, e.g., the furthest encoded source (LF3) has the largest NRMSE ($5.57$). In addition, we observe in Figure 9 that while the general trends are the same across the two embeddings (e.g., LF3 and LF1 are the furthest and closest to HF), probabilistic embedding more accurately quantifies model form uncertainties especially in the case of highly biased LF sources.
Figure 10 compares the prediction accuracy on unseen data with probabilistic and deterministic embeddings. As it can be observed, the probabilistic approach is slightly more robust in HF emulation but both approaches $(1)$ provide the same degree of accuracy and robustness for LF sources, and $(2)$ are less accurate in emulating HF and LF3 sources. The reason LF is learnt less accurately than other LF sources is its low correlation while the errors in emulating the HF source primarily stem from the lack of HF data.
@@ 5.3 Inverse Parameter Estimation
In this section, we compare the performance of GP+ in inverse parameter estimation against the method of KOH and the open-source software package UQLab [106]. As schematically demonstrated in Figure 3(a), KOH uses an additive bias and can calibrate a single LF source at a time. In our studies we consider two version of this method where its parameters (e.g., the kernel parameters of LF and HF sources) are estimated either jointly or via a modular scheme [28, 86, 107] which first estimates the parameters of the LF source and then optimizes the rest of the parameters. The calibration module of UQLab relies on Bayesian inference and leverages MCMC for parameter estimation. Similar to previous sections, we repeat each of our studies $10$ times and report the average values.
@@
We now use the high-dimensional Borehole problem presented in Table 5 with $1$ HF and $2$ LF sources that both have nonlinear model form errors (note that LF1 is more biased compared to LF2, see the NRMSEs in Table 5). We generate $20$ and $100$ samples from, respectively, the HF and each of the LF sources. We only corrupt the HF training samples with noise and use $1800$ noise-free HF samples for evaluation of models’ performance. There are two calibration parameters in this example and their ground truth values are $\zeta_{1}=250$ and $\zeta_{2}=1500$. We compare the calibration results obtained by GP+ with those from KOH whose parameters are estimated either jointly or by a modular approach. Since KOH can only fuse two sources at a time, we consider different combinations of LF and HF sources in our experiments.
Figure 11: High-fidelity emulation performance in Borehole: GP+ consistently outperforms KOH’s approach. The superior performance of both variations of GP+ are primarily attributed to using all the data jointly and dispensing with the assumption that model form error are additive.
@@
The superiority of GP+ in emulation is coupled with a more accurate estimation of the calibration parameters as illustrated in Figure 12. While the estimated values are quite accurate across all methods, probabilistic GP+ provides slightly better results, especially compared to modular KOH that fuses the data in a sequence of steps rather than jointly. We also observe that estimations for $\zeta_{1}$ in Figure 11(a) are generally more accurate than those of $\zeta_{2}$ in Figure 11(b). This trend is primarily due to the fact that the underlying functions (i.e., HF and both LF sources) are more sensitive to $\zeta_{1}$ than $\zeta_{2}$ (see Table 6 for sensitivity analysis of Borehole).
@@
Figure 13 compares the performance of GP+ in predicting unseen HF data against both the modular and joint variations of KOH’s method. As it can be observed, the probabilistic GP+ achieves the lowest NRMSE as well as NIS and is closely followed by its deterministic counterpart. As demonstrated in Figure 14 we observe a similar trend in estimating the calibration parameters where both version of GP+ not only provide estimates that are closer to the ground truth values, but also show more robustness to variations in the training data. Comparing the estimates across the four calibration parameters, we notice that all the models are less accurate in the case of $\zeta_{1}$. We attribute this behavior to identifiability issues and the fact that the underlying functions are almost insensitive to $\zeta_{1}$ (see the sensitivity analysis in Table 8).
@@ 5.4 Bayesian Optimization
BO is a global optimization method that is increasingly used in optimizing black-box and expensive-to-evaluate objective functions. The two main ingredients of BO are an acquisition function (AF) [108, 109, 110, 111, 112, 113] and an emulator which iteratively interact while searching for the global optimum. In BO, GPs are dominantly used for emulation [114, 115, 116, 117, 118, 119] since they are easy and fast to train, can effectively learn from small data, and naturally provide prediction uncertainties which are needed in the AF.
Given the widespread use of GPs in BO, we equip GP+ with a few unique functionalities that streamline optimization of black-box and expensive-to-evaluate objective functions. As schematically demonstrated in Figure 15, GP+ enables MFBO with just a few lines of codes; primarily with the BO function which has a few features that distinguish it from other BO packages such as BoTorch. Firstly, it leverages the emulator described in Sections 4.1, 4.2 and 4.3 which provides more accuracy than competing GP modeling packages. Secondly, it has the option to tailor the emulation process to BO by integrating MAP with scoring rules. As detailed in Appendix G, this integration improves the accuracy of GPs’ prediction intervals and, in turn, improves the exploration aspect of BO in the context of MF problems [87]. Lastly, the BO function in GP+ has a specialized AF that quantifies the information value of HF and LF data such that they are used primarily for exploitation and exploration, respectively [3]. At iteration $k$ during MFBO, this AF quantifies the value of a sample as:
@@
The results of our comparison studies are summarized in Figure 16 where the top row illustrates the fidelity embeddings learned for each example by GP+. These embeddings are learnt based on the initial MF data and provide a quantitative metric for assessing the relative accuracy of each LF source with respect to the HF source. Based on these embeddings, while all the fidelity sources are globally correlated in HEA, two of the LF sources in HOIP and Borehole are highly biased and have a limited potential to improve optimization. The effects of these correlated/uncorrelated sources on the optimization are illustrated in the bottom row of Figure 16 where the convergence histories are provided by tracking the best HF estimate found by each method (i.e., $y_{l}^{*}$ in Section 5.4) as a function of the accumulated sampling cost. Specifically, as shown in Figure 15(d), the inexpensive correlated LF sources of HEA significantly improve the BO where MFBO finds the same compound as SFBOEI but at a much lower cost. We note that both MFBO and SFBOEI converge before finding the smallest HF value primarily because both the HF and LF data are inherently noisy.
@@
(d) HEA
(e) Borehole
(f) HOIP
Figure 16: Fidelity manifolds and convergence histories in BO problems: The top row displays the fidelity embedding that is learned by GP+ based on the initial data in each example. These embeddings quantify the global correlation among the LF sources and the HF source. The bottom row shows the convergence history of each method versus their accumulated cost. The thin curves illustrate each repetition and the thick curves indicate the average behavior across the $10$ repetitions. Based on these plots, GP+ significantly outperforms SFBOEI when LF sources are sufficiently correlated with the HF source. However, in scenarios without any correlated sources (HOIP), GP+ may converge prematurely due to its second stop condition that limits the number of iterations.
As shown in Figure 15(e), the superior performance of MFBO is more evident in the Borehole example which has two informative LF sources (LF3 and LF4 based on Figure 15(b)). In this case, MFBO is able to effectively leverage the LF sources in exploring the input space and occasionally samples from the HF source even though the sampling cost associated with it is very high. This infrequent sampling reduces the overall cost but is necessary for converging to the true minimum. Unlike MFBO, BoTorch fails in this example since $(1)$ its MF emulator is inaccurate, and $(2)$ its AF incorrectly quantifies the information value to the extent that it cannot find an HF candidate whose value warrants its high sampling cost. As expected, SFBOEI converges to a value that is quite close to the minimum but at a much higher cost than MFBO.
HOIP is an example where the interpretable diagnostic tools of GP+ prove useful. Specifically, the fidelity embedding in Figure 15(c) indicates that both LF sources are highly biased compared to the HF one and hence optimizing the latter may not benefit from sampling from the LF sources especially if the cost ratios are unbalanced. By investigating the optimization histories in Figure 15(f) we realize that MFBO is unable to provide the same improvements as in the other two examples. Specifically, MFBO finds an optimum that is quite close to the one found by SFBOEI ($-15.87$ vs $-16.2$) but it does so at a much lower cost ($2000$ vs $6000$) since it primarily samples from the LF sources, see Figure 17. The performance of MFBO can be improved in such applications with highly biased LF sources by initializing the optimization process with more HF data.
@@ 6 Conclusions and Future Directions
In this paper we primarily focused on applications where data (especially high-fidelity data) is scarce and hence probabilistic modeling has an edge over a deterministic one. We achieve this probabilistic nature in GP+ via variational approaches and plan to extend it in our future works based on MCMC. We anticipate this extension to more accurately quantify uncertainties but with a noticeable increase in computational costs. Another interesting future direction is integrating our contributions with techniques that extend GPs to solve partial differential equations, classification tasks, or big data problems.
A unique advantage of GP+ over other GP modeling libraries is the interpretability of its learnt latent spaces. One example is the fidelity embedding that GP+ constructs for MF data which provides a visualizable metric that quantifies the relative similarity of the data sources. Based on the formulations in Sections 4.1 and 4.2, this fidelity embedding quantifies the global correlations among the data sources since it is not a function of the input features $\boldsymbol{x}$ or $\boldsymbol{t}$. We plan to extend this approach to learn local correlations but note that this extension will rely on larger training datasets since the required embedding functions will have more parameters. With a somewhat similar approach, we believe the calibration scheme introduced in Section 4.4 can be generalized to cases where not all the unknown parameters are shared across the computer models. This generalization is tightly connected to the so-called non-identifiability issues which are both worthy of in-depth future investigations.

### v2

@@ 5.2 Multi-fidelity Modeling
We use two analytic (Sinusoidal and Wing) and one engineering (DNS-ROM) examples for the comparison (see Table 4 and [deng2023data] for details on these examples). Sinusoidal is a $1$-dimensional, bi-fidelity example for which we generate a dataset consisting of $400$ HF and $2000$ LF samples. Wing has $4$ fidelity sources ($1$ HF and $3$ LFs) and we produce $1500$ samples from the HF source and $4000$ samples from each of the LF sources. In both analytic examples, $1\%$ of data is used for training and the rest of the HF data for testing. DNS-ROM is a $5$-dimensional problem on fracture modeling of metallic alloys where the data are generated via four different simulators with $70,110,170,250$ samples. In this example, we use $20\%$ of the samples for training and the rest of the HF data for testing.
The results for each approach on each problem are summarized in Table 2 and demonstrate that GP+ significantly outperforms the other baselines in all problems. More specifically, while V-GP is limited to the small HF data, GP+ effectively leverages the information provided by the LF data to learn the HF source. The poor performance of STMF-GP is due to the fact that from a methodological standpoint it models the inter-relations between the data sources incorrectly. In addition to providing low accuracy, the pblackictions of STMF-GP are sensitive to the values assigned to its fidelity indices. This is evident in Table 2 where the pblackiction errors for two different yet close sets of random indices ($STMF-GP_{1}$ and $STMF-GP_{2}$, see also Table 11) are very different. Regarding FFNNs, we attribute their poor performance in all problems to their architecture and, in particular, their simple mechanism for handling fidelity levels. These FFNNs simply one-hot encode the fidelity indices and ignore the intricate correlations among the corresponding data sources. Compablack to other methods, the reported NRMSEs for FFNNs are more sensitive to the model architecture and notably change as the network size varies. This sensitivity is partly due to the small size of the MF data and can perhaps be improved by iteratively refining the architecture or the optimization parameters (e.g., learning rate schedule or regularization weights). However, we avoid such detailed refinements since none of the other methods are fine-tuned.
@@
Wing and DNS-ROM are relatively complex problems with small amounts of data and different types of noise (e.g., inDNS-ROM the noise variance depends on the source while in Wing it does not). GP+ is very well suited to tackle these types of problems because the number of its hyperparameters scales much better than FFNNs, is not limited to the small HF data, and better estimates noise as explained in Section 4.1.1. Accordingly, we observe lower pblackiction errors for GP+ in these examples.
@@ 5.2.1 Deterministic and Probabilistic Embedding
Figure 9: Probabilistic vs deterministic embedding for MF modeling: While both embeddings estimate the same degree of similarity among sources, the probabilistic one characterizes more uncertainties.
@@
Figure 10: Pblackiction performance with probabilistic vs deterministic embedding: We report NRMSE and NIS values calculated across $10$ repetitions. Both approaches are very accurate but the probabilistic one is more robust for HF emulation especially since HF data is very small. The similar NRMS and NIS values for LF1 and LF2 are expected because of their high correlation.
Figure 9 illustrates the fidelity embeddings learned through probabilistic and deterministic MF modeling. We only show the learnt embeddings in one of the repetitions since the relative distances across the $10$ repetitions are quite similar (albeit the exact locations are different). As explained in Section 4, these embeddings indicate how similar or correlated different data sources are with respect to each other. Specifically, the latent distance between the points encoding the LF sources from the point that encodes the HF source are consistent with the NRMSE values reported in Table 4 where, e.g., the furthest encoded source (LF3) has the largest NRMSE ($5.57$). In addition, we observe in Figure 9 that while the general trends are the same across the two embeddings (e.g., LF3 and LF1 are the furthest and closest to HF), probabilistic embedding more accurately quantifies model form uncertainties especially in the case of highly biased LF sources.
Figure 10 compares the pblackiction accuracy on unseen data with probabilistic and deterministic embeddings. As it can be observed, the probabilistic approach is slightly more robust in HF emulation but both approaches $(1)$ provide the same degree of accuracy and robustness for LF sources, and $(2)$ are less accurate in emulating HF and LF3 sources. The reason LF is learnt less accurately than other LF sources is its low correlation while the errors in emulating the HF source primarily stem from the lack of HF data.
@@ 5.3 Inverse Parameter Estimation
In this section, we compare the performance of GP+ in inverse parameter estimation against the method of KOH and the open-source software package UQLab [marelli2014uqlab]. As schematically demonstrated in Figure 3(a), KOH uses an additive bias and can calibrate a single LF source at a time. In our studies, we consider two versions of this method where its parameters (e.g., the kernel parameters of LF and HF sources) are estimated either jointly or via a modular scheme [RN791, RN270, RN645] which first estimates the parameters of the LF source and then optimizes the rest of the parameters. The calibration module of UQLab relies on Bayesian inference and leverages MCMC for parameter estimation. Similar to previous sections, we repeat each of our studies $10$ times and report the average values.
@@
Figure 11: High-fidelity emulation performance in Borehole: GP+ consistently outperforms KOH’s approach. The superior performance of both variations of GP+ are primarily attributed to using all the data jointly and dispensing with the assumption that model form errors are additive.
We now use the high-dimensional Borehole problem presented in Table 5 with $1$ HF and $2$ LF sources that both have nonlinear model form errors (note that LF1 is more biased compablack to LF2, see the NRMSEs in Table 5). We generate $20$ and $100$ samples from, respectively, the HF and each of the LF sources. We only corrupt the HF training samples with noise and use $1800$ noise-free HF samples for evaluation of models’ performance. There are two calibration parameters in this example and their ground truth values are $\zeta_{1}=250$ and $\zeta_{2}=1500$. We compare the calibration results obtained by GP+ with those from KOH whose parameters are estimated either jointly or by a modular approach. Since KOH can only fuse two sources at a time, we consider different combinations of LF and HF sources in our experiments.
@@
The superiority of GP+ in emulation is coupled with a more accurate estimation of the calibration parameters as illustrated in Figure 12. While the estimated values are quite accurate across all methods, probabilistic GP+ provides slightly better results, especially compablack to modular KOH that fuses the data in a sequence of steps rather than jointly. We also observe that estimations for $\zeta_{1}$ in Figure 12(a) are generally more accurate than those of $\zeta_{2}$ in Figure 12(b). This trend is primarily due to the fact that the underlying functions (i.e., HF and both LF sources) are more sensitive to $\zeta_{1}$ than $\zeta_{2}$ (see Table 6 for sensitivity analysis of Borehole).
@@
Figure 13 compares the performance of GP+ in pblackicting unseen HF data against both the modular and joint variations of KOH’s method. As it can be observed, the probabilistic GP+ achieves the lowest NRMSE as well as NIS and is closely followed by its deterministic counterpart. As demonstrated in Figure 14 we observe a similar trend in estimating the calibration parameters where both versions of GP+ not only provide estimates that are closer to the ground truth values, but also show more robustness to variations in the training data. Comparing the estimates across the four calibration parameters, we notice that all the models are less accurate in the case of $\zeta_{1}$. We attribute this behavior to identifiability issues and the fact that the underlying functions are almost insensitive to $\zeta_{1}$ (see the sensitivity analysis in Table 8).
@@ 5.4 Bayesian Optimization
BO is a global optimization method that is increasingly used in optimizing black-box and expensive-to-evaluate objective functions. The two main ingblackients of BO are an acquisition function (AF) [turner2021bayesian, frazier2015bayesian, couckuyt2022bayesian, nguyen2023tutorial, brochu2010tutorial, foumani2023multi] and an emulator which iteratively interact while searching for the global optimum. In BO, GPs are dominantly used for emulation [kopsiaftis2019gaussian, binois2022survey, sui2018stagewise, morita2022applying, mcintire2016sparse, rana2017high] since they are easy and fast to train, can effectively learn from small data, and naturally provide pblackiction uncertainties which are needed in the AF.
Given the widespread use of GPs in BO, we equip GP+ with a few unique functionalities that streamline optimization of black-box and expensive-to-evaluate objective functions. As schematically demonstrated in Figure 15, GP+ enables MFBO with just a few lines of codes; primarily with the BO function which has a few features that distinguish it from other BO packages such as BoTorch. First, it leverages the emulator described in Sections 4.1, 4.2 and 4.3 which provides more accuracy than competing GP modeling packages. Second, it has the option to tailor the emulation process to BO by integrating MAP with scoring rules. As detailed in Appendix G, this integration improves the accuracy of GPs’ pblackiction intervals and, in turn, improves the exploration aspect of BO in the context of MF problems [foumani2023effects]. Lastly, the BO function in GP+ has a specialized AF that quantifies the information value of HF and LF data such that they are used primarily for exploitation and exploration, respectively [RN1845]. At iteration $k$ during MFBO, this AF quantifies the value of a sample as:
@@
The results of our comparison studies are summarized in Figures 16 and 17 where Figure 16 illustrates the fidelity embeddings learned for each example by GP+. These embeddings are learnt based on the initial MF data and provide a quantitative metric for assessing the relative accuracy of each LF source with respect to the HF source. Based on these embeddings, while all the fidelity sources are globally correlated in HEA, two of the LF sources in HOIP and Borehole are highly biased and have a limited potential to improve optimization. The effects of these correlated/uncorrelated sources on the optimization are illustrated in Figure 17 where the convergence histories are provided by tracking the best HF estimate found by each method (i.e., $y_{l}^{*}$ in Section 5.4) as a function of the accumulated sampling cost. Specifically, as shown in Figure 17(a), the inexpensive correlated LF sources of HEA significantly improve the BO where MFBO finds the same compound as SFBOEI but at a much lower cost. We note that both MFBO and SFBOEI converge before finding the smallest HF value primarily because both the HF and LF data are inherently noisy.
@@
Figure 16: Fidelity manifolds in BO problems: The figures display the fidelity embeddings learned by GP+ based on the initial data in each example. These embeddings quantify the global correlation among the LF sources and the HF source. Based on these plots, while the LF sources in Borehole and HEA  are sufficiently correlated with the HF source, both LF sources of HOIP reveal low correlation with the HF source. This can cause premature convergence of GP+ due to its second stop condition that limits the number of iterations.
(a) HEA
(b) Borehole
(c) HOIP
Figure 17: Convergence histories in BO problems: The figures show the convergence history of each method versus their accumulated cost. The thin curves illustrate each repetition and the thick curves indicate the average behavior across the $10$ repetitions. Based on these plots, GP+ significantly outperforms SFBOEI when LF sources are sufficiently correlated with the HF source (see also Figure 16(a)). However, in scenarios without any correlated sources (HOIP, see Figure 16(c)), GP+ may converge prematurely due to its second stop condition that limits the number of iterations.
As shown in Figure 17(b), the superior performance of MFBO is more evident in the Borehole example which has two informative LF sources (LF3 and LF4 based on Figure 16(b)). In this case, MFBO is able to effectively leverage the LF sources in exploring the input space and occasionally samples from the HF source even though the sampling cost associated with it is very high. This infrequent sampling blackuces the overall cost but is necessary for converging to the true minimum. Unlike MFBO, BoTorch fails in this example since $(1)$ its MF emulator is inaccurate, and $(2)$ its AF incorrectly quantifies the information value to the extent that it cannot find an HF candidate whose value warrants its high sampling cost. As expected, SFBOEI converges to a value that is quite close to the minimum but at a much higher cost than MFBO.
HOIP is an example where the interpretable diagnostic tools of GP+ prove useful. Specifically, the fidelity embedding in Figure 16(c) indicates that both LF sources are highly biased compablack to the HF one and hence optimizing the latter may not benefit from sampling from the LF sources especially if the cost ratios are unbalanced. By investigating the optimization histories in Figure 17(c) we realize that MFBO is unable to provide the same improvements as in the other two examples. Specifically, MFBO finds an optimum that is quite close to the one found by SFBOEI ($-15.87$ vs $-16.2$) but it does so at a much lower cost ($2000$ vs $6000$) since it primarily samples from the LF sources, see Figure 18. The performance of MFBO can be improved in such applications with highly biased LF sources by initializing the optimization process with more HF data.
@@ 6 Conclusions and Future Directions
In this paper, we primarily focused on applications where data (especially high-fidelity data) is scarce and hence probabilistic modeling has an edge over a deterministic one. We achieve this probabilistic nature in GP+ via variational approaches and plan to extend it in our future works based on MCMC. We anticipate this extension to more accurately quantify uncertainties but with a noticeable increase in computational costs. Additionally, all the problems analyzed in this paper were single-output. Currently, GP+ manages multi-output problems by concatenating the input space with categorical features to distinguish among the outputs. We aim to add more options for handling multi-output problems in future versions of GP+. Another interesting future direction is integrating our contributions with techniques that extend GPs to solve partial differential equations, classification tasks, or big data problems.
A unique advantage of GP+ over other GP modeling libraries is the interpretability of its learnt latent spaces. One example is the fidelity embedding that GP+ constructs for MF data which provides a visualizable metric that quantifies the relative similarity of the data sources. Based on the formulations in Sections 4.1 and 4.2, this fidelity embedding quantifies the global correlations among the data sources since it is not a function of the input features $\boldsymbol{x}$ or $\boldsymbol{t}$. We plan to extend this approach to learn local correlations but note that this extension will rely on larger training datasets since the requiblack embedding functions will have more parameters. With a somewhat similar approach, we believe the calibration scheme introduced in Section 4.4 can be generalized to cases where not all the unknown parameters are shablack across the computer models. This generalization is tightly connected to the so-called non-identifiability issues which are both worthy of in-depth future investigations.